# High-Speed Event Vision-Based Tactile Roller Sensor for Large Surface Measurements

### v1

@@ III-F Braille Feature Recognition
Without motion compensation, if we were to simply create a frame-like image by plotting all events recorded within a specific time window at their original pixel locations (a process called direct accumulation), the relative movement between the sensor and the Braille plate would cause significant motion blur. This occurs because events generated by a single point on the moving Braille surface would be spread across multiple pixels in the resulting image (as seen in Figure 6a). Using a long accumulation time integrates more events but exacerbates this blur as features traverse many pixels, whereas a very short accumulation time minimizes blur but often captures too few events to form a clear representation, leading to degraded features [10, 57]. To counteract this, we warp events with batches consisting of a constant number of events ($N_{e}$ = 20,000 events) at the initial timestamp of the event stream as the reference time, $t_{ref}$. For each event $e_{k}=(\mathbf{x}_{k},t_{k},p_{k})$ where $\mathbf{x}_{k}=(x_{k},y_{k})$, its warped position $\mathbf{x}^{\prime}_{k}$ is calculated as:
@@
From one such generated frame $I_{warp}$, we select a Region of Interest (ROI) corresponding to a single Braille cell location. This ROI, illustrated in Figure 6b, then undergoes an image processing pipeline to enhance dot visibility and reduce noise. The steps include: applying a bilateral filter for edge-preserving smoothing, using Contrast Limited Adaptive Histogram Equalization (CLAHE) to increase local contrast, performing binary thresholding to segment potential dots, and applying morphological opening and closing operations to eliminate salt-and-pepper noise.
Figure 6: Braille feature recognition pipeline applied on single letter ROI. (a) Direct accumulation of events (b) Motion compensated image of letter. (b) Denoised and processed letter image. (c) Hough transform applied and corresponding letter identified.
IV Experimental Setup
@@
  • Several 3D printed PLA objects with diverse geometries (e.g., stairs with fine gradations, sparse sphere (concentric circles) - see Fig. 8a.) were used to test the reconstruction capabilities on varied shapes.
  • A larger 3D printed line array with varying heights (Fig. 8b) was employed to evaluate the sensor’s performance in large surface reconstruction tasks involving stitching multiple frames.
@@
The subsequent Results section (Sec. V) details the outcomes of these experiments. The evaluations focus on: (1) 3D reconstruction accuracy, primarily quantified using Mean Absolute Error (MAE) against ground truth geometries, including an analysis of the proposed BMA fusion strategy; (2) the relationship between rolling speed and reconstruction accuracy; (3) the effectiveness of large surface reconstruction through measurement stitching; and (4) the speed and accuracy achieved during high-speed Braille character recognition.
V Results
(a) Refer to caption
(b) Refer to caption
Figure 8: Examples of 3D surface reconstruction results using the neuromorphic roller sensor. (a) Single object reconstruction (top view): 4mm sphere, 0.1mm stairs, sparse sphere. (b) Large surface reconstruction: 3D printed varying-height line array (top and isometric views).
V-A Single Object Reconstruction
@@
As seen in Table III, the BMA fusion framework outperforms EMVS with an average percentage decrease in MAE of 25.2% across all objects tested. Additionally, the fusion strategy of only fusing the left and right images that performed the best in reconstruction variability with Li et al. [53] was tested. While their method performs better than using the depth map generated using baseline EMVS, the proposed BMA fusion framework still outperforms the other methods. The depth maps generated for the single object reconstruction experiments in this comparison are shown in Figure 8a. Also, a qualitative analysis of the fusion strategy was conducted. A line array (3 identical rectangles with height 0.6 mm) was reconstructed using EMVS and BMA fusion framework. The pointclouds, depicted side-by-side in Figure 9, showcase a greater depth consistency and lower curvature warping when BMA is utilized. This visual improvement in consistency is confirmed quantitatively; the standard deviation of the reconstructed depth across the line array decreased significantly from 0.067 mm for the baseline EMVS reconstruction to 0.028 mm when using the BMA fusion approach, indicating a much more uniform depth estimation.
@@
TABLE III: Evaluation of the tactile 3D reconstruction Average MAE in mm using different depth map fusion weights at 0.3 m/s
Reconstruction Method Stairs Sphere Sparse Sphere Average
@@
An analysis of the reconstruction accuracy across different object geometries (Table III) reveals a consistent performance trend: the sparse sphere yielded the lowest MAE, followed by the stairs, with the solid sphere exhibiting the highest MAE. This trend correlates strongly with the operating principle of event cameras, which primarily generate signals at intensity edges during relative motion under consistent illumination. The sparse sphere presents distinct, well-separated edges with significant depth variance, facilitating robust event generation and accurate reconstruction. The stairs object, conversely, features numerous edges with smaller depth variations (0.1–0.2 mm) in close proximity, potentially leading to denser but possibly less distinct event clusters and increased reconstruction error. Finally, the solid sphere, possessing a smooth, continuous geometry without sharp edges (aside from the contact boundary), generates fewer internal feature-related events compared to the other objects (as suggested by the sparse sphere reconstruction in Fig. 8a), resulting in a sparser representation and consequently the highest reconstruction MAE among the tested shapes.
Figure 9: Comparison of line array point cloud reconstruction for EMVS and BMA respectively. BMA pointcloud exhibits less noisy events and less curvature produced from cylindrical roller. Moreover, the MAE is lower by 25.2% when BMA fusion is used, highlighting decrease in depth inconsistency and curvature errors. Color represents change in depth using JET colormap.
@@
BMA also generated depth maps with lower MAE across varying speeds as illustrated in Figure 10 in comparison to EMVS and Left & Right Fusion. However, the reduction in the MAE of the sphere is more significant at higher speeds, with a 5.27% reduction at 0.05 m/s, a 25.2% reduction at 0.3 m/s and starts to plateau with a 26.4% reduction at 0.5 m/s in comparison to EMVS. Interestingly, Figure 10 also shows that at lower speeds (e.g., 0.05 m/s), the performance difference between the BMA fusion (using all three reference views) and the Left & Right fusion (using only start and end views) becomes negligible. This suggests that while BMA fusion provides substantial benefits at higher speeds by effectively combining information from more distinct viewpoints to mitigate motion artifacts and improve consistency, at lower speeds the depth maps generated at the start, middle, and end times are likely very similar due to the minimal camera displacement within the time window. Consequently, fusing only the start and end views captures most of the available information, and adding the highly correlated midpoint view offers little additional advantage. These findings are in line with Bhattacharya et al. [58] as their event-based depth estimation model exhibits greater depth prediction quality at higher speeds. This effect could be explained by the fact that more events are generated within the same time window at higher speeds allowing for more accurate depth estimation through EMVS.
@@
V-B Large Surface Reconstruction
@@
The large surface reconstruction capabilities of our roller were tested with a 3D printed line array with varying heights (0.1-1.2 mm at intervals of 0.1 mm). The final point cloud is generated by aligning results from multiple depth maps as done by Zhang et al. [49]. The multiple depth maps are aligned using ICP; since the speed and hence the translation of the roller are known from the robotic controller, the translation along the rolling direction is used as the initial pose for ICP. The reconstruction results can be seen in Figure 8b with a resulting MAE of 0.0638 mm across the whole surface. As aforementioned, speed is critical for large surface reconstruction using tactile sensing. Our roller exhibits, to the best of our knowledge, the highest measured speed at 0.5 m/s or 500 mm/s for tactile scanning with an MAE below 100 microns. This is 11 times faster than the maximum reported speed for rolling belt tactile sensors (45 mm/s) [26] and 45 times faster than the highest speed for roller tactile sensors (11 mm/s) [24].
@@
Table IV summarizes the performance trade-offs for different continuous tactile sensing approaches based on the provided documents. Our roller significantly advances the state-of-the-art in terms of speed, achieving 500 mm/s while maintaining a competitive 3D reconstruction accuracy (54.9 $\mu m$ MAE). GelBelt demonstrates the potential for very high accuracy ($\sim$13.5 $\mu m$ average MAE, measured against a high-resolution GelSight Max scan as ground truth for specific defects) but operates at a considerably lower maximum speed of 45 mm/s [26]. The GelRoller paper focused on a novel self-supervised reconstruction method, achieving 175 µm MAE in single-frame tests, but did not report a maximum scanning speed [49]. TouchRoller operated at the slowest speed (11 mm/s) and was designed for 2D texture mapping, not 3D reconstruction [24].
@@
TABLE IV: Comparison between Tactile Belt/Roller Sensor Performance
@@
Ours 500 54.9
V-C High-Speed Tactile Braille Reading

### v2

@@ IV-D Braille Feature Recognition
Without motion compensation, if we were to simply create a frame-like image by plotting all events recorded within a specific time window at their original pixel locations (a process called direct accumulation), the relative movement between the sensor and the Braille plate would cause significant motion blur. This occurs because events generated by a single point on the moving Braille surface would be spread across multiple pixels in the resulting image (as seen in Figure 12a). Using a long accumulation time integrates more events but exacerbates this blur as features traverse many pixels, whereas a very short accumulation time minimizes blur but often captures too few events to form a clear representation, leading to degraded features [Salah2023, Gallego_CMAX_2018]. To counteract this, we warp events with batches consisting of a constant number of events ($N_{e}$ = 20,000 events) at the initial timestamp of the event stream as the reference time, $t_{ref}$. For each event $e_{k}=(\mathbf{x}_{k},t_{k},p_{k})$ where $\mathbf{x}_{k}=(x_{k},y_{k})$, its warped position $\mathbf{x}^{\prime}_{k}$ is calculated as:
@@
V Experimental Setup And Procedures
V-A Experimental Setup
@@
  • Several 3D printed PLA objects with diverse geometries: stairs with fine gradations, sparse sphere (concentric circles), hexagon with a shallow recess, bumps or semi-cylinders side by side with varying radii, and a pyramid (see Fig. 8) were used to test the reconstruction capabilities on varied shapes.
  • Two larger 3D printed plates were used to evaluate performance in large surface reconstruction tasks involving stitching multiple frames (see Fig. 11): a line array with varying heights and a letter plate with varying-height letters and induced defects (e.g., scratches and printing artifacts).
@@
The subsequent Results section (Sec. VI) details the outcomes of these experiments. The evaluations focus on: (1) 3D reconstruction accuracy, primarily quantified using Mean Absolute Error (MAE) against ground truth geometries, including an analysis of the proposed BMA fusion strategy; (2) the relationship between rolling speed and reconstruction accuracy; (3) the effectiveness of large surface reconstruction through measurement stitching; and (4) the speed and accuracy achieved during high-speed Braille character recognition. The MAE is calculated by comparing the reconstructed depth map, $Z_{f}$, directly to the ground truth depth map, $Z_{gt}$. The ground truth map is created by first identifying the contact patch in the reconstructed data and then using the known physical geometry of the object to compute the ideal depth value for each pixel within that masked patch. This procedure ensures that $Z_{gt}$ is inherently pixel-wise aligned with $Z_{f}$. The MAE is then the average of the absolute difference in depth values between the two maps. All reported MAE values are computed only on the sparse measurement points output by our method. Pixels without a valid depth estimate (shown as white in the visualized point clouds in Section VI) are excluded from the error computation.
V-B Calibration Cross-Validation and Evaluation
A potential concern from the calibration performed, as described in Section IV-C, is that the optimal fusion weights determined using only the sphere geometry might be biased and not generalize well to other object shapes (outlined in detail in Section V). To investigate this, we performed a cross-validation analysis. For calibration, the sensor is rolled over the surface of the calibration sphere/object (rather than pressing the sphere into a static roller as is done typically with VBTS/NVBTS) while recording event data. From each rolling pass, a point cloud is generated and the optimization in Eq. 8 is performed within the masked contact region to determine the optimal parameters. We first determined the object-specific optimal fusion parameters $\theta_{i}$ (the weights $w_{s},w_{m},w_{e}$ were highlighted since the thresholding parameters did not change across objects) by running the grid search optimization (Eq. 8) using the ground truth $Z_{gt}^{(i)}$ for each object $i$ (where $i\in\{\text{Sphere, Stairs, Sparse Sphere}\}$).
We then evaluated how well each set of parameters $\theta_{i}$ performed on average across all $M=3$ test objects. For each object $i$, the calibration procedure was repeated for five independent rolling trials to ensure consistency of the estimated parameters. The MAE for object $i$ was first computed for each trial and then averaged over the five trials to obtain a single representative MAE value for that object. Using these per-object averages, we then calculated the overall average MAE, $e_{i}$, for parameter set $\theta_{i}$ across all $M=3$ objects as follows:
$e_{i}=\frac{1}{M}\sum_{i=1}^{M}\bigl\|Z_{f}^{(i)}(\theta_{i})-Z_{gt}^{(i)}\bigr\|_{1}$ (11)
where $Z_{f}^{(j)}(\theta_{i})$ is the fused depth map generated for object $i$ using parameters $\theta_{i}$. The values reported in Table II are therefore the result of a two-stage averaging process: first averaging the MAE over the five trials for each object, and then averaging these per-object MAEs across the three objects for each parameter set. We then compared these error values to determine which calibration object yielded the parameters with the best overall generalization performance (i.e., the minimum $e_{i}$).
The results are summarized in Table II. While the object-specific optimal weights vary slightly, the weights calibrated using the sphere yielded the lowest overall average MAE ($e_{\text{Sphere}}=0.0549$ mm) when applied to the objects within the calibration set, confirming its suitability as a general calibration target.
TABLE II: Cross-validation of BMA fusion weights
Calibrated on $w_{s}$ $w_{m}$ $w_{e}$ Avg. MAE of all objects (mm)
Sphere 0.445 0.333 0.222 0.0549
Stairs 0.5 0.333 0.167 0.0652
Sparse Sphere 0.4 0.4 0.2 0.0649
A consistent and perhaps counter-intuitive trend observed across all calibrations is that the start reference view ($t_{s}$) receives the highest (or tied for highest) weight, while the end view ($t_{e}$) consistently receives the lowest. Intuitively, one might expect the midpoint view ($t_{m}$), representing a temporal average, to be most reliable. However, this weighting suggests the information captured early in the time window is most critical for accurate reconstruction in our setup. The possible reasons for this trend and its implications are discussed in more detail in Section VII.
V-C Contact Force Analysis
To evaluate the effect of normal contact force on reconstruction quality, we conducted a qualitative analysis. Using the UR10 robot’s precise position control, we varied the vertical distance between the roller and the tactile plate to apply three distinct levels of elastomer deformation, corresponding to light, medium, and maximum contact forces. Figure 7 illustrates the resulting motion-compensated event frames generated from the bumps object under these conditions.
The results show a clear relationship between contact force and signal quality. A light pressing force results in minimal deformation with the elastomer barely touching the plate, leading to weak event generation where critical features of the object such as the smaller semi-cylinders are not captured. In contrast, a maximum pressing force ensures the elastomer fully conforms to the surface geometry. This produces a strong, clear signal with high event counts, capturing the full geometry of the object and thus enabling the most accurate reconstruction. Based on this analysis, all results presented in this paper were obtained using a maximum contact force setting to ensure consistency and optimal data quality.
Figure 7: Qualitative analysis of the effect of pressing force on signal quality using the bumps object. Motion-compensated event frames generated under light, medium, and maximum contact forces (left to right).
VI Results
VI-A Single Object Reconstruction
@@
As seen in Table III, the BMA fusion framework outperforms EMVS with an average percentage decrease in MAE of 25.2% across objects tested within the calibration set. Additionally, the fusion strategy of only fusing the left and right images that performed the best in reconstruction variability with Li et al. [li_2025_cyclic] was tested. While their method performs better than using the depth map generated using baseline EMVS, the proposed BMA fusion framework still outperforms the other methods. A quantitative analysis on a line array (three identical rectangles with height 0.6 mm) further confirms the improvement in depth consistency: the standard deviation of the reconstructed depth across the line array decreased significantly from 0.067 mm for the baseline EMVS reconstruction to 0.028 mm when using the BMA fusion approach, indicating a much more uniform depth estimation.
@@
TABLE III: Evaluation of the tactile 3D reconstruction Average MAE in $m$ using different depth map fusion weights at 0.3 m/s with calibration set
Reconstruction Method Stairs (mm) Sphere (mm) Sparse Sphere (mm) Average (mm)
@@
The comprehensive reconstruction accuracy for all six tested objects, each averaged over five independent trials, is presented in Table IV. An analysis of these results reveals a consistent performance trend that correlates strongly with the operating principle of event cameras, which primarily generate signals at intensity edges during relative motion under consistent illumination. The sparse sphere yielded the lowest MAE, as its distinct, well-separated edges with significant depth variance facilitate robust event generation. The stairs and pyramid objects, both possess well-defined, repeating features, but with smaller depth variations and in close proximity. These features could have potentially led to denser but less distinct event clusters and increased reconstruction error compared to the sparse sphere, although they were also reconstructed with high accuracy. The bumps object presents a hybrid case; the heights of the semi-cylinders and the sections between them were captured accurately, but their smooth, curved tops were not reconstructed. This is consistent with the limitations of the event-based algorithm on curved surfaces that lack edges or sharp features, similar to the solid sphere. The hexagon and sphere objects exhibited the highest MAE respectively. With regards to the hexagon, as seen in Figure 8, the reconstruction of the hexagon is incomplete, failing to capture the entire central recess. This can be attributed to self-occlusion, where the steep walls of the deformed elastomer block the camera’s line of sight or perhaps the elastomer’s Shore Hardness was not suitable to fully conform to the recess and capture its features accurately, leading to higher MAE overall. The solid sphere, possessing a smooth, continuous geometry without sharp edges (aside from the contact boundary), generates fewer internal feature-related events compared to the other objects (Fig. 8), resulting in a sparser representation and consequently the highest reconstruction MAE among the tested shapes.
Figure 8: Single object reconstruction. Objects from top to bottom: 4mm sphere, stairs, sparse sphere, hexagon with recess, bumps, pyramid.
@@
BMA also generated depth maps with lower MAE across varying speeds as illustrated in Figure 9 in comparison to EMVS and Left & Right Fusion. However, the reduction in the MAE of the sphere is more significant at higher speeds, with a 5.27% reduction at 0.05 m/s, a 25.2% reduction at 0.3 m/s and starts to plateau with a 26.4% reduction at 0.5 m/s in comparison to EMVS. Interestingly, Figure 9 also shows that at lower speeds (e.g., 0.05 m/s), the performance difference between the BMA fusion (using all three reference views) and the Left & Right fusion (using only start and end views) becomes negligible. This suggests that while BMA fusion provides substantial benefits at higher speeds by effectively combining information from more distinct viewpoints to mitigate motion artifacts and improve consistency, at lower speeds the depth maps generated at the start, middle, and end times are likely very similar due to the minimal camera displacement within the time window. Consequently, fusing only the start and end views captures most of the available information, and adding the highly correlated midpoint view offers little additional advantage. These findings are in line with Bhattacharya et al. [bhattacharya2024monocular] as their event-based depth estimation model exhibits greater depth prediction quality at higher speeds. This effect could be explained by the fact that more events are generated within the same time window at higher speeds allowing for more accurate depth estimation through EMVS.
Another important consideration when it comes to BMA is the optimal number of reference views for fusion. To investigate this, we conducted an additional experiment on the sphere object using five reference views. This approach yielded a higher MAE of 0.1115 ± 0.0065 mm compared to the 0.0946 ± 0.0081 mm from the three-reference fusion, albeit with lower variability. This suggests a trade-off where adding more views can increase robustness but may also introduce additional warping errors that degrade overall accuracy. This finding is consistent with observations by Li et al. [li_2025_cyclic], who noted that excessive fusion can weaken unique information from individual viewpoints. The optimal number of reference views is likely task-dependent and differs depending on speed used, but for the conditions in our experiments, three views provided the best balance of accuracy and consistency.
TABLE IV: Reconstruction Accuracy for All Single Objects
Object Average MAE (mm) Std. Dev. (mm) Sparse Sphere 0.0120 0.0010 Stairs 0.0581 0.0026 Bumps 0.0589 0.0024 Pyramid 0.0677 0.0058 Hexagon 0.0808 0.0042 Sphere 0.0946 0.0081 Overall 0.0620 0.0047
@@
To further assess the sensor’s qualitative performance and generalization, we tested it on three additional objects with diverse materials and surface complexities, as shown in Figure 10. The scanned objects include a rivet head with a surface defect, an M2 screw with fine threads, and a textured fabric. The resulting point clouds in each case exhibit strong geometric consistency with the physical objects. These results demonstrate the sensor’s robustness in reconstructing a variety of complex geometries beyond the tested 3D printed shapes and potential for industrial surface inspection applications.
Figure 10: Qualitative reconstruction results. (a) Physical objects: a rivet head with a defect, an M2 screw, and a textured fabric. (b) Corresponding 3D point cloud reconstructions. Color represents change in depth using JET colormap
VI-B Large Surface Reconstruction
Figure 11: Examples of large surface reconstruction. (a) 3D printed varying-height line array (top) and its corresponding reconstruction (bottom). (b) 3D printed letter plate with varying-height letters and induced defects (top) and its reconstruction (bottom).
The large surface reconstruction capabilities of our roller were tested with two 3D printed plates, shown in Figure 11. The final point cloud for each is generated by aligning results from multiple depth maps as done by Zhang et al. [zhang_gelroller_2024]. The multiple depth maps are aligned using ICP; since the speed and hence the translation of the roller are known from the robotic controller, the translation along the rolling direction is used as the initial pose for ICP. It is important to note that due to the high repeatability of the UR10 robot and the short travel distances in our experiments, the initial pose derived from proprioception is already highly accurate. While the ICP step provides minimal refinement in these specific tests, it is retained as a standard practice consistent with prior work [zhang_gelroller_2024] to ensure the pipeline’s robustness and generalizability for larger-area scans where robot drift could become a factor. The robot’s motion error within a single 20 ms reconstruction window is negligible compared to other error sources and does not significantly impact the precision of an individual depth map.
The test objects’ reconstruction accuracy was averaged over five trials as was done for the single object reconstruction. The first test object was a line array with varying heights (0.1-1.2 mm at intervals of 0.1 mm), for which our sensor achieved a resulting MAE of 0.0638 $\pm$ 0.0055 mm across the whole surface (see Figure 11a). The second object was a letter plate, designed to benchmark against prior work in letter reconstruction like GelRoller [zhang_gelroller_2024] and to test defect detection (see Figure 11b). The letters on this plate have small, varying heights (ranging from 0.4 mm to 0.6 mm), and the plate includes both intentionally modeled and induced defects like scratches and printing artifacts. For the letter plate, our sensor achieved an MAE of 0.0596 ± 0.0035 mm, and as shown qualitatively in Figure 11b, it successfully captured the letter geometries. Also, the fine defects on the surface were also captured.
@@ VI-B Large Surface Reconstruction
As aforementioned, speed is critical for large surface reconstruction using tactile sensing. Our roller exhibits, to the best of our knowledge, the highest measured speed at 0.5 m/s or 500 mm/s for tactile scanning with an MAE below 100 microns. This is 11 times faster than the maximum reported speed for rolling belt tactile sensors (45 mm/s) [mirzaee_gelbelt_2025] and 45 times faster than the highest speed for roller tactile sensors (11 mm/s) [Cao2023].
@@
Table V summarizes the performance trade-offs for different continuous tactile sensing approaches based on the provided documents. Our roller significantly advances the state-of-the-art in terms of speed, achieving 500 mm/s while maintaining a competitive 3D reconstruction accuracy (62.0 $\mu m$ MAE). GelBelt demonstrates the potential for very high accuracy ($\sim$13.5 $\mu m$ average MAE, measured against a high-resolution GelSight Max scan as ground truth for specific defects) but operates at a considerably lower maximum speed of 45 mm/s [mirzaee_gelbelt_2025]. The GelRoller paper focused on a novel self-supervised reconstruction method, achieving 175 µm MAE in single-frame tests, but did not report a maximum scanning speed [zhang_gelroller_2024]. TouchRoller operated at the slowest speed (11 mm/s) and was designed for 2D texture mapping, not 3D reconstruction [Cao2023].
@@
TABLE V: Comparison between Tactile Belt/Roller Sensor Performance
@@
Ours 500 62.0
VI-C High-Speed Tactile Braille Reading
The large surface tactile scanning of the roller was also tested with regards to feature detection and recognition. Two braille plates were scanned with standard braille sizing: semi sphere of radius of 0.635 mm and spacing of 2.54 mm. To obtain the Braille character recognition results, we processed the generated IWE, $I_{warp}$, (from Section IV-D) with an application-specific pipeline. From $I_{warp}$, we select a Region of Interest (ROI) corresponding to a single Braille cell location. This ROI, illustrated in Figure 12b, then undergoes an image processing pipeline to enhance dot visibility and reduce noise. The steps include: applying a bilateral filter for edge-preserving smoothing, using Contrast Limited Adaptive Histogram Equalization (CLAHE) to increase local contrast, performing binary thresholding to segment potential dots, and applying morphological opening and closing operations to eliminate salt-and-pepper noise.
Figure 12: Braille feature recognition pipeline applied on single letter ROI. (a) Direct accumulation of events (b) Motion compensated image of letter. (c) Denoised and processed letter image. (d) Hough transform applied and corresponding letter identified.